# Measurement incompatibility in Bayesian multiparameter quantum estimation

### v1

@@ I Introduction
Quantum metrology explores the fundamental and practical impact of quantum mechanical phenomena, such as superposition, squeezing, and entanglement, on high-precision measurements of physical parameters [giovannetti2011advances, Polino2020, Liu2022c, Jiao2023a, Jiao2023a, Montenegro2025b]. Its central aim is to surpass the precision limits of classical strategies that do not rely on these quantum effects. Even if such advantage is not always feasible, a proper understanding of the fundamental precision limits imposed by quantum mechanics is crucial to develop new technologies, guiding how to optimally allocate resources [demkowiczdobrzanski2015quantum, sahota2015quantum, friis2017flexible].
@@
The mathematical tools to analyse the metrological precision achievable with quantum systems are the subject of quantum estimation theory (QET), its most celebrated result being the quantum Cramér–Rao bound (CRB) [helstrom1976quantuum]. While many results in quantum metrology fundamentally rely on the CRB, this approach has important limitations. Roughly speaking, the CRB is a meaningful figure of merit only when a large number of experiment repetitions can be performed, with a significant caveat that it is often hard to understand a priori how many repetitions are needed. This flavour of QET is known as local, meaning that its predictions are generally valid when the discrepancy between the estimates and the true value of the parameters is small enough, which eventually happens as more and more samples are accumulated.
@@
Global estimation theory represents a complementary approach that is especially relevant when resources are constrained and only a limited number of experimental trials can be performed [tsang2012zivzakai, rubio2019quantum, valeri2020experimental]. While several alternative approaches have been proposed [Gessner2023, Maccone2017, Gebhart2024, Hervas2025, Hervas2025a, Meyer2025, Mukhopadhyay2025a, Polloreno2023, Allen2025], the most established and extensively developed framework for going beyond the local estimation paradigm is Bayesian quantum estimation. Dating back to the early days of the field [personick1969thesis, personick1971application, Holevo1973, Holevo1973b, yuen1973multipleparameter, helstrom1974noncommuting, helstrom1976quantuum], it is nowadays being increasingly applied to a wide range of multiparameter sensing experiments involving both photons [DAurelio2022a, Valeri2023, Cimini2024, Cimini2024a, Belliardo2024b, Minati2025] and atoms [glatthard2022optimal, overton2024five]. This is the framework adopted in the present work.
@@
Bayesian estimation is not only practically significant in regimes with limited experimental data but also holds deep fundamental implications. For example, it clarifies impractical claims arising from applying local QET beyond its valid domain, such as the alleged effectiveness of super-Heisenberg-limited schemes [rivas2012subheisenberg, tsang2012zivzakai, pezze2013subheisenberg] or the apparent enhancement in multiple phase estimation [humphreys2013quantum, knott2016local, rubio2020bayesian, Gorecki2022], and it reveals the need to correct the Heisenberg limit in phase estimation by a factor of $\pi$ [gorecki2020pi]. In general, the strategies predicted by local QET can become suboptimal in the non-asymptotic regime [Meyer2025], where the Bayesian framework consistently enables efficient information extraction [rubio2019quantum, ravell2025knowledge]. This is unsurprising given that Bayesian inference constitutes the most basic form of learning [jaynes2003probability], which is the most comprehensive approach for characterising quantum systems [Gebhart2023a, Berquin2023, Banchi2024a, Chen2024h, Oh2024, Dutt2024, Wang2025c, Hu2025a, Belliardo2025].
@@
One of the fundamental aspects of quantum theory is the existence of incompatible measurements that cannot be implemented jointly [Heinosaari2016, Guhne2023]. In the context of quantum metrology, this phenomenon becomes crucial when multiple parameters must be estimated simultaneously: measurement incompatibility can impose fundamental trade-offs on the achievable precision of different parameters [Szczykulska2016, Li2024x, Hu2025f, Kull2020], which have been quantified in various ways [Razavian2020, Belliardo2021, Candeloro2021c, Candeloro2024]. Besides this concrete consequence, incompatibility makes multiparameter theory very challenging from a mathematical point of view [demkowicz2020multi]. While the fundamental precision limits in single-parameter quantum metrology are relatively well-understood, fundamental multiparameter limits remain a subject of active research [Albarelli2019c, Liu2019d, Pezze2025].
@@
It is instructive to briefly review the impact of such incompatibility effects on the development of local QET, before drawing a parallel with the Bayesian approach. A multiparameter version of the quantum CRB was first derived by Helstrom in the dawn of the field, based on the notion of the symmetric logarithmic derivative (SLD) [Helstrom1967, helstrom1976quantuum]. We shall call this bound the Helstrom (SLD) CRB. While this bound proved attainable for single-parameter estimation [Braunstein1994, Barndorff-Nielsen2000], it does not account for the effects of incompatibility and is not generally attainable for multiparameter estimation. In the following decades, further quantum generalisations of the CRB were developed, providing a better account of measurement incompatibility, and culminating in the discovery that the ultimate precision limit attainable with collective (entangling) measurements over many copies is given by the Holevo–Cramér–Rao bound (HCRB) [demkowicz2020multi, Hayashi2008a, Kahn2009, Yamagata2013, Yang2018a, Fujiwara2023]. It was eventually understood that, in this asymptotic scenario where many copies can be measured collectively, the effect of measurement incompatibility is relatively mild: the HCRB can be at most twice the original bound obtained by Helstrom [Tsang2019, Carollo2019] (see also Refs. [Albarelli2019c, demkowicz2020multi, Belliardo2021]).
@@
By contrast, the fundamental precision achievable with only separable measurements on each copy of the state is not efficiently computable in local QET, even numerically (although collective measurements are unnecessary for pure states [Matsumoto2002]). Nonetheless, several bounds for this setting have been derived [Gill2000, Lu2020a, Chen2022f, Chen2022g, Chen2024j, Wang2025f], among which we highlight the Nagaoka–Hayashi (NH) bound [Nagaoka1989, Nagaoka1991, hayashi1999simultaneous, Conlon2020, Conlon2024c, Yung2024, Das2024, Conlon2025a]. Most recently, the problem has been studied using conic programming [Hayashi2023a], providing a unifying framework and new tools that have already been applied to probe state optimisation [Hayashi2024].
@@
Despite significant advances in local QET, the role of measurement incompatibility in Bayesian estimation has received much less attention. The general conditions that a Bayesian estimation strategy must satisfy to be optimal were already identified in the aforementioned early works [Holevo1973, Holevo1973b, helstrom1974noncommuting, helstrom1976quantuum], including a discrete analogue for hypothesis testing [Yuen1970, Yuen1975]. A generalisation to convex functions has also been proposed [Coutts2018]. helstrom1976quantuum provided examples of how these conditions can be used to verify whether a proposed strategy is indeed optimal. Reversing this procedure—namely, using the conditions to derive the optimal strategy—is, however, exceedingly challenging, with only a limited number of problems where this can be done analytically. For instance, this procedure can be carried out in the simple case of two hypotheses [brody1996bayesian].
@@
For single-parameter estimation, the fairly broad class of location-isomorphic estimation problems can be optimised exactly [rubio2024first] and the solution can be shown to satisfy the corresponding optimality conditions. This family includes, as a particular case, the optimisation of the standard mean square error—called mean square loss (MSL) in this work—a problem originally solved by Personick [personick1971application]. While Personick made an early attempt to extend this reasoning to the estimation of two parameters [personick1969thesis], closed-form solutions based on indirect measurements were subsequently obtained [Belavkin1973], and a solution to the problem of estimating two displacement parameters encoded in Gaussian states with Gaussian priors is known [wang2007], a general and practical procedure for determining optimal solutions in the Bayesian estimation of multiple parameters has remained elusive for decades.
@@
Only recently has a multiparameter bound, analogous to the Helstrom (SLD) CRB—and thus neglecting incompatibility effects—been derived [rubio2020bayesian]. As motivated later, we propose referring to this as the symmetric posterior mean (SPM) bound. Bayesian analogues of several other local bounds, including the NH bound, have since been established, incorporating measurement incompatibility to a certain extent [Suzuki2024a, Zhang2024g]. In addition, numerical algorithms for identifying optimal solutions are now under active investigation [Bavaresco2024designing].
@@
In this manuscript, we show that, in analogy with local QET, the minimum MSL can be at most twice the SPM bound obtained by neglecting measurement incompatibility. Unlike the local QET result this holds on single-shot, however, we also find that this upper bound can be trivial, i.e., larger than the MSL based only on a priori knowledge. This arises from the central role of prior information in Bayesian estimation, which, as demonstrated through examples in quantum imaging, phase and dephasing estimation, and qubit sensing, can influence the perceived incompatibility of the estimation protocol. For this reason, we further derive a tighter upper bound by exploiting ideas from hypothesis testing [Belavkin1975a, Belavkin1975, Kholevo1979, Hausladen1994], and employ lower bounds tighter than the SPM bound, particularly the Bayesian NH bound [Suzuki2024a]. Furthermore, we provide open-source code for the numerical evaluation of all relevant bounds, available on GitHub [multipar2025code]. Finally, the manuscript aims to serve as a pedagogical and self-contained introduction to Bayesian multiparameter quantum metrology, offering a foundation for further research in the field.
@@ II.1 Preliminaries
For simplicity, in Sec. IV, we only consider examples with discrete outcomes, which is natural when focusing on finite-dimensional systems. In such cases, continuous outcomes can be realised by performing finite measurements depending on a classical parameter [chiribella2007how]. Conversely, discrete outcomes can be embedded in a continuous-outcome formulation by expressing $M(x)$ as a sum of Dirac delta functions centered on the discrete outcome $n$,
@@
associated with a POVM element $M_{n}$ satisfying $M_{n}\succeq 0$ and $\sum_{n}M_{n}=\mathbb{I}$. The continuous-outcome formalism is particularly convenient for variational calculus, which is frequently employed in Bayesian optimisation [personick1971application, jaynes2003probability].
@@
This is the so-called maximum a posteriori (MAP) estimator [linden2014bayesian], which plays a role similar to that of the maximum likelihood estimator in local QET [kay1993fundamentals].
@@ II.2 Estimation performance
We first observe that Eq. (7) quantifies estimation performance by integrating over all hypothesis values in $\Omega$, whose contribution can vary from small to large, weighted by the prior probability $p(\boldsymbol{\theta})$. This is the key feature that makes the framework not only global but also Bayesian. There exists a more conservative, non-Bayesian approach to global estimation: the minimax approach, which is based on the frequentist loss [demkowiczdobrzanski2015quantum]
@@
and seeks to minimise the worst-case value of $\mathcal{L}(\boldsymbol{\theta})$ over relevant regions of the parameter space [gorecki2023heisenberg]. However, since the minimax criterion can be written as the supremum of Eq. (7) over all priors, our bounds in Sec. III are also relevant for minimax estimation. Note that, at the other extreme, Eq. (8) is employed in local QET by evaluating it in an infinitesimal neighbourhood of specific values of $\boldsymbol{\theta}$ [kay1993fundamentals, demkowiczdobrzanski2015quantum].
We further comment that, in Bayesian estimation, the additional integral over $x$ is required only when searching for optimal estimation protocols, as in our case, since this task precedes the recording of specific outcomes. For data post-processing, estimation performance must instead be assessed using an outcome-dependent quantifier [jaynes1976confidence], e.g.,
@@
Alternatively, one may employ credibility intervals, also based on the posterior [jaynes1976confidence, boeyens2021uninformed]. Both experimental error assessment and the minimax and local estimation approaches discussed above are beyond the scope of this work, and we therefore focus exclusively on Eq. (7).
@@
where $L$ is a positive semidefinite weight matrix, i.e., $L\succeq 0$, and we assume it to be parameter-independent. The resulting $\mathcal{L}$ is the standard MSL, and this loss quantifier leads to the posterior mean (PM) as the optimal Bayes estimator [jaynes2003probability]:
@@
There is a widespread belief that the MSL applies universally, regardless of the specific system, measurement, or type of parameters of interest. Strictly speaking, however, the MSL is only appropriate for location or shift parameters. This is because the topology for such parameters corresponds to the real line, so that the corresponding loss function must be translation invariant [jaynes2003probability], i.e.,
@@
In contrast, parameter types with different properties require alternative loss functions. Cycle parameters, for instance, are estimated using trigonometric loss functions—phases in interferometry are the archetypical example [helstrom1976quantuum, holevo2011probabilistic, demkowiczdobrzanski2015quantum]. Moreover, recent results show that scale parameters such as temperature or decay rate require logarithmic loss functions for consistent estimation [rubio2021nov], while weight parameters such as probability of success or coherence parameters demand hyperbolic loss functions [rubio2025on]. Naive use of the MSL, while it may occasionally yield correct results—particularly when $\boldsymbol{\Theta}$ is known to lie close to a reference point $\boldsymbol{\Theta}_{0}$—is physically unjustified and therefore discouraged unless there is no better choice.
Despite this, one can retain the mathematical simplicity of the MSL while respecting parameters’ more general topologies in certain scenarios. Specifically, Refs. [rubio2024first, tsang2022generalized] demonstrated that many parameter types admit mean losses equivalent to the MSL, up to a mapping that, in the context of multiparameter estimation, can be expressed as
@@
where $\boldsymbol{f}$ is chosen according to symmetries of the problem [rubio2025on]; see Ref. [overton2024five] for an experimental demonstration with cold $\mathrm{Cs}$ atoms. Note that phases fall outside this framework [holevo2011probabilistic], though they can be approximated by the MSL itself when confined to intervals of order $\pi/2$ [friis2017flexible, rubio2019quantum].

### v2

@@ I Introduction
Quantum metrology explores the fundamental and practical impact of quantum mechanical phenomena, such as superposition, squeezing, and entanglement, on high-precision measurements of physical parameters [1, 2, 3, 4, 4, 5]. Its central aim is to surpass the precision limits of classical strategies that do not rely on these quantum effects. Even if such advantage is not always feasible, a proper understanding of the fundamental precision limits imposed by quantum mechanics is crucial to develop new technologies, guiding how to optimally allocate resources [6, 7, 8].
@@
The mathematical tools to analyse the metrological precision achievable with quantum systems are the subject of quantum estimation theory (QET), its most celebrated result being the quantum Cramér–Rao bound (CRB) [9]. While many results in quantum metrology fundamentally rely on the CRB, this approach has important limitations. Roughly speaking, the CRB is a meaningful figure of merit only when a large number of experiment repetitions can be performed, with a significant caveat that it is often hard to understand a priori how many repetitions are needed. This flavour of QET is known as local, meaning that its predictions are generally valid when the discrepancy between the estimates and the true value of the parameters is small enough, which eventually happens as more and more samples are accumulated.
@@
Global estimation theory represents a complementary approach that is especially relevant when resources are constrained and only a limited number of experimental trials can be performed [10, 11, 12]. While several alternative approaches have been proposed [13, 14, 15, 16, 17, 18, 19, 20, 21], the most established and extensively developed framework for going beyond the local estimation paradigm is Bayesian quantum estimation. Dating back to the early days of the field [22, 23, 24, 25, 26, 27, 9], it is nowadays being increasingly applied to a wide range of multiparameter sensing experiments involving both photons [28, 29, 30, 31, 32, 33] and atoms [34, 35]. This is the framework adopted in the present work.
@@
Bayesian estimation is not only practically significant in regimes with limited experimental data but also holds deep fundamental implications. For example, it clarifies impractical claims arising from applying local QET beyond its valid domain, such as the alleged effectiveness of super-Heisenberg-limited schemes [36, 10, 37] or the apparent enhancement in multiple phase estimation [38, 39, 40, 41], and it reveals the need to correct the Heisenberg limit in phase estimation by a factor of $\pi$ [42]. In general, the strategies predicted by local QET can become suboptimal in the non-asymptotic regime [18], where the Bayesian framework consistently enables efficient information extraction [11, 43]. This is unsurprising given that Bayesian inference constitutes the most basic form of learning [44], which is the most comprehensive approach for characterising quantum systems [45, 46, 47, 48, 49, 50, 51, 52, 53].
@@
One of the fundamental aspects of quantum theory is the existence of incompatible measurements that cannot be implemented jointly [54, 55]. In the context of quantum metrology, this phenomenon becomes crucial when multiple parameters must be estimated simultaneously: measurement incompatibility can impose fundamental trade-offs on the achievable precision of different parameters [56, 57, 58, 59], which have been quantified in various ways [60, 61, 62, 63]. Besides this concrete consequence, incompatibility makes multiparameter theory very challenging from a mathematical point of view [64]. While the fundamental precision limits in single-parameter quantum metrology are relatively well-understood, fundamental multiparameter limits remain a subject of active research [65, 66, 67].
@@
It is instructive to briefly review the impact of such incompatibility effects on the development of local QET, before drawing a parallel with the Bayesian approach. A multiparameter version of the quantum CRB was first derived by Helstrom in the dawn of the field, based on the notion of the symmetric logarithmic derivative (SLD) [68, 9]. We shall call this bound the Helstrom (SLD) CRB. While this bound proved attainable for single-parameter estimation [69, 70], it does not account for the effects of incompatibility and is not generally attainable for multiparameter estimation. In the following decades, further quantum generalisations of the CRB were developed, providing a better account of measurement incompatibility, and culminating in the discovery that the ultimate precision limit attainable with collective (entangling) measurements over many copies is given by the Holevo–Cramér–Rao bound (HCRB) [64, 71, 72, 73, 74, 75]. It was eventually understood that, in this asymptotic scenario where many copies can be measured collectively, the effect of measurement incompatibility is relatively mild: the HCRB can be at most twice the original bound obtained by Helstrom [76, 77] (see also Refs. [65, 64, 61]).
@@
By contrast, the fundamental precision achievable with only separable measurements on each copy of the state is not efficiently computable in local QET, even numerically (although collective measurements are unnecessary for pure states [78]). Nonetheless, several bounds for this setting have been derived [79, 80, 81, 82, 83, 84], among which we highlight the Nagaoka–Hayashi (NH) bound [85, 86, 87, 88, 89, 90, 91, 92]. Most recently, the problem has been studied using conic programming [93], providing a unifying framework and new tools that have already been applied to probe state optimisation [94].
@@
Despite significant advances in local QET, the role of measurement incompatibility in Bayesian estimation has received much less attention. The general conditions that a Bayesian estimation strategy must satisfy to be optimal were already identified in the aforementioned early works [24, 25, 27, 9], including a discrete analogue for hypothesis testing [95, 96]. A generalisation to convex functions has also been proposed [97]. Helstrom [9] provided examples of how these conditions can be used to verify whether a proposed strategy is indeed optimal. Reversing this procedure—namely, using the conditions to derive the optimal strategy—is, however, exceedingly challenging, with only a limited number of problems where this can be done analytically. For instance, this procedure can be carried out in the simple case of two hypotheses [98].
@@
For single-parameter estimation, the fairly broad class of location-isomorphic estimation problems can be optimised exactly [99] and the solution can be shown to satisfy the corresponding optimality conditions. This family includes, as a particular case, the optimisation of the standard mean square error—called mean square loss (MSL) in this work—a problem originally solved by Personick [23]. While Personick made an early attempt to extend this reasoning to the estimation of two parameters [22], closed-form solutions based on indirect measurements were subsequently obtained [100], and a solution to the problem of estimating two displacement parameters encoded in Gaussian states with Gaussian priors is known [101], a general and practical procedure for determining optimal solutions in the Bayesian estimation of multiple parameters has remained elusive for decades.
@@
Only recently has a multiparameter bound, analogous to the Helstrom (SLD) CRB—and thus neglecting incompatibility effects—been derived [40]. As motivated later, we propose referring to this as the symmetric posterior mean (SPM) bound. Bayesian analogues of several other local bounds, including the NH bound, have since been established, incorporating measurement incompatibility to a certain extent [102, 103]. In addition, numerical algorithms for identifying optimal solutions are now under active investigation [104, 105].
@@
Contributions.—We show that, in analogy with local QET, the minimum MSL can be at most twice the SPM benchmark that is obtained when measurement incompatibility is ignored. Key differences, however, are that this upper bound now holds at the single-shot level, but can also be trivial, i.e., larger than the MSL derived solely from a priori information. This effect stems from the central role of prior information in Bayesian estimation, which, as illustrated through examples in quantum imaging, phase and dephasing estimation, and qubit sensing, may significantly modify the apparent incompatibility of the estimation protocol. Consequently, measurement incompatibility in Bayesian multiparameter quantum estimation is not only a relatively mild factor in the overall estimation precision, but it can also be concealed when the influence of the prior is strong. We further derive a tighter upper bound that is guaranteed to be non-trivial, by combining measurements introduced in hypothesis testing [106, 107, 108, 109] with optimal classical Bayes estimators [44]. Our upper bounds are complementary to existing lower bounds tighter than the SPM one, where the Bayesian NH bound [102] represents the state of the art. In addition, we derive a class of lower bounds based on generalised posterior mean operators beyond the symmetric version that, while less tight than the NH bound, may be of independent interest, since they are obtained via the theory of monotone metrics, analogously to established results in local QET [110]. An open-source code for the numerical evaluation of all relevant bounds is available on GitHub [111]. Finally, beyond deriving bounds on the minimum MSL, we determine explicit conditions for the measurement strategies that achieve it. Combined with the pedagogical and self-contained introduction to Bayesian multiparameter quantum metrology presented in this work, these results establish a solid foundation for further advances in the field.
@@ II.1 Preliminaries
For simplicity, in Sec. IV, we only consider examples with discrete outcomes, which is natural when focusing on finite-dimensional systems. In such cases, continuous outcomes can be realised by performing finite measurements depending on a classical parameter [112]. Conversely, discrete outcomes can be embedded in a continuous-outcome formulation by expressing $M(x)$ as a sum of Dirac delta functions centered on the discrete outcome $n$,
@@
associated with a POVM element $M_{n}$ satisfying $M_{n}\succeq 0$ and $\sum_{n}M_{n}=\mathbb{I}$. The continuous-outcome formalism is particularly convenient for variational calculus, which is frequently employed in Bayesian optimisation [23, 44].
@@
This is the so-called maximum a posteriori (MAP) estimator [113], which plays a role similar to that of the maximum likelihood estimator in local QET [114].
@@ II.2 Estimation performance
We first observe that Eq. (7) quantifies estimation performance by integrating over all hypothesis values in $\Omega$, whose contribution can vary from small to large, weighted by the prior probability $p(\boldsymbol{\theta})$. This is the key feature that makes the framework not only global but also Bayesian. There exists a more conservative, non-Bayesian approach to global estimation: the minimax approach, which is based on the frequentist loss [6]
@@
and seeks to minimise the worst-case value of $\mathcal{L}(\boldsymbol{\theta})$ over relevant regions of the parameter space [115]. However, since the minimax criterion can be written as the supremum of Eq. (7) over all priors, our bounds in Sec. III are also relevant for minimax estimation. Note that, at the other extreme, Eq. (8) is employed in local QET by evaluating it in an infinitesimal neighbourhood of specific values of $\boldsymbol{\theta}$ [114, 6].
We further comment that, in Bayesian estimation, the additional integral over $x$ is required only when searching for optimal estimation protocols, as in our case, since this task precedes the recording of specific outcomes. For data post-processing, estimation performance must instead be assessed using an outcome-dependent quantifier [116], e.g.,
@@
Alternatively, one may employ credibility intervals, also based on the posterior [116, 117]. Both experimental error assessment and the minimax and local estimation approaches discussed above are beyond the scope of this work, and we therefore focus exclusively on Eq. (7).
@@
where $L$ is a positive semidefinite weight matrix, i.e., $L\succeq 0$, and we assume it to be parameter-independent. The resulting $\mathcal{L}$ is the standard MSL, and this loss quantifier leads to the posterior mean (PM) as the optimal Bayes estimator [44]:
@@
There is a widespread belief that the MSL applies universally, regardless of the specific system, measurement, or type of parameters of interest. Strictly speaking, however, the MSL is only appropriate for location or shift parameters. This is because the topology for such parameters corresponds to the real line, so that the corresponding loss function must be translation invariant [44], i.e.,
@@
In contrast, parameter types with different properties require alternative loss functions. Cycle parameters, for instance, are estimated using trigonometric loss functions—phases in interferometry are the archetypical example [9, 118, 6]. Moreover, recent results show that scale parameters such as temperature or decay rate require logarithmic loss functions for consistent estimation [119], while weight parameters such as probability of success or coherence parameters demand hyperbolic loss functions [120]. Naive use of the MSL, while it may occasionally yield correct results—particularly when $\boldsymbol{\Theta}$ is known to lie close to a reference point $\boldsymbol{\Theta}_{0}$—is physically unjustified and therefore discouraged unless there is no better choice.
Despite this, one can retain the mathematical simplicity of the MSL while respecting parameters’ more general topologies in certain scenarios. Specifically, Refs. [99, 121] demonstrated that many parameter types admit mean losses equivalent to the MSL, up to a mapping that, in the context of multiparameter estimation, can be expressed as
@@
where $\boldsymbol{f}$ is chosen according to symmetries of the problem [120]; see Ref. [35] for an experimental demonstration with cold $\mathrm{Cs}$ atoms. Note that phases fall outside this framework [118], though they can be approximated by the MSL itself when confined to intervals of order $\pi/2$ [8, 11].